# Equations involving the modular $j$-function and its derivatives

### v1

@@ 7.3. Two more examples
Let us apply Theorem 1.2 to get some information on the zeroes of the function $j^{\prime\prime\prime}$. From the differential equation of the $j$-function (2.2) we see that
@@
By Theorem 1.2 (actually by Theorem 7.3), the equation
(7.2) $03j^{2}(j-1728)^{2}(j^{\prime\prime})^{2}-\left(j^{2}-1968j+2654208\right)(j^{\prime})^{4}=$ 0
has a Zariski dense set of solutions outside $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$. Then these are also solutions of $j^{\prime\prime\prime}(z)=0$.
@@
are equal to 00 at $i$ and $\rho$ and do not have exponential growth in any fundamental domain. However, we know that (7.2) has a zero $\tau\notin\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$. Therefore, the second ratio above has a pole at $\tau$, and so for all large enough $m$ the equation (7.3) has a zero near $\tau+m$. This means that (7.2), and hence $j^{\prime\prime\prime}=0$, has solutions near $-\frac{1}{\tau+m}$ which accumulate at 00.
@@
Given an equation $F(z,\mathbf{j}(z))=0$, our strategy of the proof of Theorem 1.2 is to apply a generic $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-transformation and show that in the function $F(\gamma z,\mathbf{j}(\gamma z))$ the ratio of the coefficients of the lowest and highest powers of $z$ has a pole at some point in $\tau$ or has exponential growth at a cusp. In some cases, e.g. when $F(X,\mathbf{Y})$ does not depend on $X$, we can keep things simple and just use the good old transformation $z\mapsto-\frac{1}{z}$. Indeed, in this case it turns out that the coefficient of the lowest power of $z$ does not depend on $j^{\prime\prime}$ and so we can apply Corollary 6.8. For instance,
@@
We give an example to show that the transformation $-\frac{1}{z}$ and even $a-\frac{1}{z}$ for any integer $a$ does not suffice in general (when $F$ depends on $X$), as the aforementioned coefficient may depend on $j^{\prime\prime}$. Consider the function $f(z)=2j^{\prime}(z)+zj^{\prime\prime}(z)$. After a $z\mapsto a-\frac{1}{z}$ transformation we get

### v3

@@ 7.3. Two more examples
Let us apply Theorem˜1.2 to get some information on the zeroes of the function $j^{\prime\prime\prime}$. From the differential equation of the $j$-function (2.2) we see that
@@
By Theorem˜1.2, the equation
(7.2) $3j^{2}(j-1728)^{2}(j^{\prime\prime})^{2}-\left(j^{2}-1968j+2654208\right)(j^{\prime})^{4}=0$
has a Zariski dense set of solutions outside $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$ (this actually follows directly from Theorem˜7.3, because the equation is $\mathbf{j}$-homogeneous). Then these are also solutions of $j^{\prime\prime\prime}(z)=0$.
@@
are equal to 0 at $i$ and $\rho$ and do not have exponential growth in any fundamental domain. However, we know that (7.2) has a zero $\tau\notin\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$. Therefore, the second ratio above has a pole at $\tau$, and so for all large enough $m$ the equation (7.3) has a zero near $\tau+m$. This means that (7.2), and hence $j^{\prime\prime\prime}=0$, has solutions near $\frac{1}{\tau+m}$ which accumulate at 0.
@@
Given an equation $F(z,\mathbf{j}(z))=0$, our strategy of the proof of Theorem˜1.2 is to apply a generic $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-transformation and show that in the function $F(\gamma z,\mathbf{j}(\gamma z))$ the ratio of the coefficients of the lowest and highest powers of $z$ has a pole at some point in $\tau$ or has exponential growth at a cusp. In some cases, e.g. when $F(X,\mathbf{Y})$ does not depend on $X$, we can keep things simple and just use the good old transformation $z\mapsto-\frac{1}{z}$. Indeed, in this case it turns out that the coefficient of the lowest power of $z$ does not depend on $j^{\prime\prime}$ and so we can apply Corollary˜6.8. For instance,
@@
We give an example to show that the transformation $\frac{1}{z}$ and even $a-\frac{1}{z}$ for any integer $a$ does not suffice in general (when $F$ depends on $X$), as the aforementioned coefficient may depend on $j^{\prime\prime}$. Consider the function $f(z)=2j^{\prime}(z)+zj^{\prime\prime}(z)$. After a $z\mapsto a-\frac{1}{z}$ transformation we get
@@
Acknowledgement
We thank the referee for a thorough reading of the paper and for numerous suggestions that helped us improve the presentation.